OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8

OpenStudy (anonymous):

\[Given~that~\log _{\sqrt{3}}9=x-\log _{\sqrt{3}}27.Find~the~value~of~x.\]

OpenStudy (anonymous):

\[\log _{\sqrt{3}}9+\log _{\sqrt{3}}27=x\]

OpenStudy (anonymous):

must we factorize it?

OpenStudy (anonymous):

\[\log _{\sqrt{3}}\]

ganeshie8 (ganeshie8):

maybe combine using the property \[\large \log_b (x) + \log_b (y) = \log_b(x*y)\]

OpenStudy (anonymous):

\[\log _{\sqrt{3}}(9\times27)=x\]

ganeshie8 (ganeshie8):

change it to exponent form

OpenStudy (anonymous):

\[9\times27=\sqrt{3}^{x}\]

ganeshie8 (ganeshie8):

\[\large \log_b M = k\] is same as the equation \[\large M = b^k\]

ganeshie8 (ganeshie8):

Excellent !

ganeshie8 (ganeshie8):

write both sides as powers of 3

OpenStudy (anonymous):

\[3^2\times3^3=3^{\frac{ x }{ 2 }}\]

OpenStudy (anonymous):

\[2+3=\frac{ x }{ 2 }\]

OpenStudy (anonymous):

\[x=10\]

ganeshie8 (ganeshie8):

\[\large \color{red}{\checkmark}\]

ganeshie8 (ganeshie8):

good job !

OpenStudy (anonymous):

Thnx @ganeshie8 :D

ganeshie8 (ganeshie8):

yw:)